OpenStudy (anonymous):

a chemist wants to obtain 80 liters of a solution that is 12% hydrochloric acid.how many liters of 10% hydrochloric acid solution and 20% hydrochloric acid should be mixed?

OpenStudy (kropot72):

This looks familiar. Couldn't you solve the system of equations I worked out for you?

OpenStudy (anonymous):

no, i tried but i couldn't :/

OpenStudy (kropot72):

Let A = required number of liters of 10% hydrochloric acid solution. Let B = required number of liters of 20% hydrochloric acid solution. A + B = 80 .........................................(1) Number of liters of hydrochloric acid in required solution = 80 * 0.12 = 9.6 0.1A + 0.2B = 9.6 .............................(2) If we multiply both sides of equation (2) by 5 we get 0.5A + B = 48 ...................................(3) Subtracting equation (3) from equation (1) gives 0.5A = 32 ...........................................(4) Can you solve equation (4) to find A and then substitute the value of A into equation (1) to find B?

OpenStudy (anonymous):

32/0.5=64 and 80-64=16 so 10 percent one is 16 and 20percent one is 64 ?

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

the answer is wrong :/

OpenStudy (kropot72):

A is the amount of 10% solution which is 64 liters. B is the amount of 20% solution which is 16 liters.

OpenStudy (anonymous):

okay its right thank you so much again!!

OpenStudy (kropot72):

You're welcome :)